# Iwin Transformer: Hierarchical Vision Transformer using Interleaved Windows

### v1

@@ Results
ImageNet-1K. For models trained from scratch on ImageNet-1K at a resolution of 224$\times$224, Iwin Transformer variants consistently demonstrate competitive Top-1 accuracy. Iwin-T achieves 82.0%, matching ConvNeXt-T (82.1%) and surpassing Swin-T (81.3%) by 0.7%. The deeper variant, Iwin-S, reaches highest 83.4%, outperforming Swin-S (83.0%) by 0.4% and ConvNeXt-S (83.1%) by 0.3%, while maintaining similar parameters (51.6M vs. 50.0M) and FLOPs (9.0G vs. 8.7G) compared to both Swin-T and ConvNeXt-T. By increasing the embedding dimension, Iwin-B achieves 83.5%, which is close to the 83.6% of PVT-v2-B4 and falls between the 83.3% of Swin-B and the 83.8% of ConvNeXt-B. The throughput for Iwin models (e.g., Iwin-T at 729 img/s) is slightly lower than that of Swin and ConvNeXt (e.g., Swin-T at 758 img/s and ConvNeXt-T at 775 img/s).
ImageNet-22K. By leveraging large-scale ImageNet-22K pre-training, Iwin significantly enhances its performance. The pre-trained Iwin-B achieves an impressive 85.5% Top-1 accuracy at a resolution of 224$\times$224. When fine-tuned to a resolution of 384$\times$384, it reaches a robust 86.6%, surpassing Swin-B (86.4%) by 0.2% and closely aligning with the carefully configured ConvNeXt-B (86.8%). The largest variant, Iwin-L, demonstrates cutting-edge performance with 86.4% at 224$\times$224 and an impressive 87.4% at 384$\times$384, competing favorably with top-tier models like Swin-L (87.3%) and ConvNeXt-L (87.5%).
@@
This high-resolution fine-tuning capability is attributed to the collaboration of Iwin’s interleaved window attention and depthwise separable convolution, which allows the model to enjoy a global view without relying on positional encoding. This robust performance at high resolutions and the potential of collaboration to replace standard attention in models such as high-resolution image and video generation are far more important than the small gap in throughput compared to Swin.
@@
DeiT-Small/16[44] 2242 22.0 4.6 406 79.9
T2T-ViT-14[45] 2242 22.0 5.2 - 81.5
PVT-Small[46] 2242 24.5 3.8 794 79.8
Twins-S[5] 2242 24.0 2.9 979 81.7
PVT-v2-B2[47] 2242 25.4 4.0 664 82.0
PoolFormer-S36[48] 2242 31.0 5.1 764 81.4
Swin-T[11] 2242 29.0 4.5 758 81.3
ConvNeXt-T [49] 2242 29 4.5 775 82.1
@@
T2T-ViT-19[45] 2242 39.2 8.9 - 81.9
PVT-Medium[46] 2242 44.2 6.7 511 81.2
Twins-B[5] 2242 56.0 8.6 433 83.2
PVT-v2-B3[47] 2242 45.2 6.9 443 83.2
PoolFormer-M36[48] 2242 56.0 9.0 494 82.1
Swin-S[11] 2242 50.0 8.7 437 83.0
ConvNeXt-S [49] 2242 50 8.7 447 83.1
Iwin-S(ours) 2242 51.6 9.0 410 83.4
@@
DeiT-Base/16[44] 2242 86.6 17.6 273 81.8
T2T-ViT-24[45] 2242 64.1 14.1 - 82.3
PVT-Large[46] 2242 61.4 9.8 357 81.7
Twins-L[5] 2242 99.2 15.1 271 83.7
PVT-v2-B4[47] 2242 62.6 10.1 298 83.6
PoolFormer-M48[48] 2242 73.0 11.8 337 82.5
Swin-B[11] 2242 88.0 15.4 287 83.3
Swin-B[11] 3842 88.0 47.0 85 84.5
ConvNeXt-B [49] 2242 89 15.4 292 83.8
ConvNeXt-B [49] 3842 89 45.0 96 85.1
@@
R-101x3 [50] 3842 388 204.6 - 84.4
R-152x4 [50] 4802 937 840.5 - 85.4
ViT-B/16 [51] 3842 86 55.4 86 84.0
ViT-L/16 [51] 3842 307 190.7 27 85.2
Swin-B [11] 2242 88 15.4 278 85.2
Swin-B [11] 3842 88 47.0 85 86.4
ConvNeXt-B [49] 2242 89 15.4 292 85.8
ConvNeXt-B [49] 3842 89 45.1 96 86.8
@@
Swin-L [11] 2242 197 34.5 145 86.3
Swin-L [11] 3842 197 103.9 46 87.3
ConvNeXt-L [49] 2242 198 34.4 147 86.6
ConvNeXt-L [49] 3842 198 101.0 50 87.5
@@ IV-B Object Detection on COCO
We follow Swin [11] settings and evaluate the Iwin backbone on the COCO [40] for both object detection and instance segmentation tasks using Mask R-CNN [52] and Cascaded Mask R-CNN [53] with the MMDetection [54] toolbox. We use multi-scale training, AdamW optimizer and pre-trained models on ImageNet-1K.
@@ Results
Figure 6: BBox mAP and Learning Rate Progression for Iwin-T vs. Swin-T on the COCO (Cascade Mask-RCNN 3$\times$ schedule).
@@
stage 1 4$\times$ (56$\times$56) ker. 3 pad. 2 4$\times$4, 96-d, LN ker. 3 pad. 2 4$\times$4, 96-d, LN ker. 3 pad. 2 4$\times$4, 128-d, LN ker. 3 pad. 2 4$\times$4, 192-d, LN
@@ IV-C Semantic Segmentation on ADE20K
We evaluate the Iwin backbone on the ADE20K [41] semantic segmentation benchmark using UperNet [55] with the MMSegmentation [56] toolbox. The experimental settings follow Swin [11]. Training is performed for 160K iterations with a total batch size of 16 (2 images per GPU across 8 GPUs).
@@ Results
TABLE IV: Results for ADE20K semantic segmentation task. FLOPs are measured with the input size of 512$\times$2048.
@@
ResNet50[35] 28.5 183 36.7 - - -
PVTv2 B2[46] 29.1 165 45.2 - - -
ConvNeXt-T[49] - - - 60.0 939 46.0
Swin-T[11] - - - 59.9 945 44.5
@@
ResNet101[35] 47.5 260 38.8 - - -
PVTv2 B3[46] 49.0 224 47.3 - - -
ConvNeXt-S[49] - - - 82.0 1027 48.7
Swin-S[11] - - - 81.3 1038 47.6
@@
ResNeXt101-64$\times$4d[57] 86.4 - 40.2 - - -
PVTv2 B4[46] 66.3 285 47.9 - - -
ConvNeXt-B[49] - - - 122.0 1170 49.1
Swin-B[11] - - - 121.0 1188 48.1
@@ IV-D Video Recognition on Kinetics-400
In line with the Video Swin Transformer [58] settings, we propose the Video Iwin Transformer for action recognition with the MMaction2 [59] toolbox on the Kinetics-400 [42] dataset. The Video Iwin Transformer is initialized with its pre-trained Iwin model from ImageNet. Unlike the Video Swin Transformer’s intricate adaptations from 2D to 3D, which involve window shifting, masking, and relative positioning, the Iwin model only necessitates the addition of a learnable absolute position encoding in the temporal dimension, leaving other components unchanged. As illustrated in Figure 9, in Iwin 3D Attention, tokens from interleaved windows across all frames are collected for attention calculation, while depthwise separable convolution maintains its 2D operation on individual frames.
@@ Results
Figure 9: Illustration of Iwin 3D Attention. In the left image, depth-wise separable convolution is performed on each frame, as in 2D. On the right, Interleaved Window Attention is performed in 3D. This effectively approximates 3D full attention, allowing connections between any tokens within a video. The only disadvantage is that position coding is required for the time dimension. To get rid of it completely, we propose two options: perform RTR operations in the time dimension and cooperate with one-dimensional convolution; the second is to increase convolution kernel size in the time dimension to span frames. Due to paper limitations, this part will be reserved for future work.
TABLE V: Comparison on Kinetics-400. “Views” indicates # temporal clip $\times$ # spatial crop. The magnitudes are Giga ($10^{9}$) and Mega ($10^{6}$) for FLOPs and Param respectively.
@@
R(2+1)D [60] - 72.0 90.0 10 × 1 75 61.8
I3D [61] ImageNet-1K 72.1 90.3 - 108 25.0
NL I3D-101 [62] ImageNet-1K 77.7 93.3 10 × 3 359 61.8
SlowFast R101+NL [63] - 79.8 93.9 10 × 3 234 59.9
X3D-XXL [64] - 80.4 94.6 10 × 3 144 20.3
MViT-B, 32×3 [20] - 80.2 94.4 1 × 5 170 36.6
MViT-B, 64×3 [20] - 81.2 95.1 3 × 3 455 36.6
ViViT-L/16x2 [16] ImageNet-21K 80.6 94.7 4 × 3 1446 310.8
ViViT-L/16x2 320 [16] ImageNet-21K 81.3 94.7 4 × 3 3992 310.8
Swin-T [11] ImageNet-1K 78.8 93.6 4 × 3 88 28.2
Swin-S [11] ImageNet-1K 80.6 94.5 4 × 3 166 49.8
Iwin-T(ours) ImageNet-1K 79.1 93.8 4 × 3 74 29.8
Iwin-S(ours) ImageNet-1K 80.0 94.1 4 × 3 140 51.1
@@ IV-E Image Generation
We follow LightningDiT [65] settings and build a FlashDiT model to validate the effectiveness of the key component in Iwin for class-conditional image generation task on ImageNet. We replace standard self-attention with a proposed combination of interleaved window attention and depthwise separable convolution. We remove position encodings and set the convolution kernel size to $3\times 3$ and the window size to $4\times 4$, based on the latent feature map dimensions of $16\times 16$ (derived from $256\times 256$ input images through 16$\times$ downsampling), where $3\times 4=12$ closely approximates 16.
@@ Results
As shown in Table VI, our proposed FlashDiT demonstrates efficiency in image generation. It achieves competitive performance in only 56 training epoches, a fraction of the epoches required by previous state-of-the-art models such as DiT (1400) and MAR (800). Unlike these models have complexity of $32^{2}$ or $16^{2}$, FlashDiT has a computational complexity of only $253^{2}+4^{2}=$ 25 while maintaining high generation quality, achieving a gFID of 3.08 and an IS of 223.2 without classifier guidance. Therefore, FlashDiT verifies the effectiveness of the key component in Iwin as a standalone module that can seamlessly replace the self-attention module in generation models.
@@
Method Epoches Params 256$\times$256 w/o CFG 256$\times$256 w/ CFG
@@
LlamaGen [66] 300 3.1B 9.38 8.24 112.9 0.69 0.67 2.18 5.97 263.3 0.81 0.58
VAR [67] 350 2.0B - - - - - 1.80 - 365.4 0.83 0.57
MagViT-v2 [68] 1080 307M 3.65 - 200.5 - - 1.78 - 319.4 - -
MAR [69] 800 945M 2.35 - 227.8 0.79 0.62 1.55 - 303.7 0.81 0.62
@@
MaskDiT [70] 1600 675M 5.69 10.34 177.9 0.74 0.60 2.28 5.67 276.6 0.80 0.61
DiT [18] 1400 675M 9.62 6.85 121.5 0.67 0.67 2.27 4.60 278.2 0.83 0.57
SiT [71] 1400 675M 8.61 6.32 131.7 0.68 0.67 2.06 4.50 270.3 0.82 0.59
MDTv2 [72] 1080 675M - - - - - 1.58 4.52 314.7 0.79 0.65
REPA [73] 800 675M 5.90 - - - - 1.42 4.70 305.7 0.80 0.65
LightningDiT [65] 64 675M 5.14 4.22 130.2 0.76 0.62 2.11 4.16 252.3 0.81 0.58

### v2

@@ Results
ImageNet-1K. For models trained from scratch on ImageNet-1K at a resolution of $\times224$, Iwin Transformer variants consistently demonstrate competitive Top-1 accuracy. Iwin-T achieves 82.0%, matching ConvNeXt-T (82.1%) and surpassing Swin-T (81.3%) by 0.7%. The deeper variant, Iwin-S, reaches highest 83.4%, outperforming Swin-S (83.0%) by 0.4% and ConvNeXt-S (83.1%) by 0.3%, while maintaining similar parameters (51.6M vs. 50.0M) and FLOPs (9.0G vs. 8.7G) compared to both Swin-T and ConvNeXt-T. By increasing the embedding dimension, Iwin-B achieves 83.5%, which is close to the 83.6% of PVT-v2-B4 and falls between the 83.3% of Swin-B and the 83.8% of ConvNeXt-B. The throughput for Iwin models (e.g., Iwin-T at 729 img/s) is slightly lower than that of Swin and ConvNeXt (e.g., Swin-T at 758 img/s and ConvNeXt-T at 775 img/s).
ImageNet-22K. By leveraging large-scale ImageNet-22K pre-training, Iwin significantly enhances its performance. The pre-trained Iwin-B achieves an impressive 85.5% Top-1 accuracy at a resolution of $\times224$. When fine-tuned to a resolution of $\times384$, it reaches a robust 86.6%, surpassing Swin-B (86.4%) by 0.2% and closely aligning with the carefully configured ConvNeXt-B (86.8%). The largest variant, Iwin-L, demonstrates cutting-edge performance with 86.4% at $\times224$ and an impressive 87.4% at $\times384$, competing favorably with top-tier models like Swin-L (87.3%) and ConvNeXt-L (87.5%).
@@
This high-resolution fine-tuning capability is attributed to the collaboration of Iwin’s interleaved window attention and depthwise convolution, which allows the model to enjoy a global view without relying on positional encoding. This robust performance at high resolutions and the potential of collaboration to replace standard attention in models such as high-resolution image and video generation are far more important than the small gap in throughput compared to Swin.
Moreover, Iwin employs standard attention computation, which allows it to benefit from increasingly efficient attention implementations such as Flash Attention [dao2022flashattention]. In contrast, Swin’s shifted window mechanism with irregular masking cannot fully leverage these optimizations. When using PyTorch’s Scaled Dot Product Attention (SDPA) [paszke2019pytorch], Iwin nearly matches its inference throughput and even surpasses it in larger model configurations such as Iwin-L.
@@
DeiT-Small/16[touvron2020deit] 2242 22.0 4.6 406 79.9
T2T-ViT-14[T2T] 2242 22.0 5.2 - 81.5
PVT-Small[PVT_v1] 2242 24.5 3.8 794 79.8
Twins-S[chu2021twins] 2242 24.0 2.9 979 81.7
PVT-v2-B2[PVT_v2] 2242 25.4 4.0 664 82.0
PoolFormer-S36[metaformer] 2242 31.0 5.1 764 81.4
Swin-T[liu2021swin] 2242 29.0 4.5 758 81.3
ConvNeXt-T [liu2022convnet] 2242 29 4.5 775 82.1
@@
T2T-ViT-19[T2T] 2242 39.2 8.9 - 81.9
PVT-Medium[PVT_v1] 2242 44.2 6.7 511 81.2
Twins-B[chu2021twins] 2242 56.0 8.6 433 83.2
PVT-v2-B3[PVT_v2] 2242 45.2 6.9 443 83.2
PoolFormer-M36[metaformer] 2242 56.0 9.0 494 82.1
Swin-S[liu2021swin] 2242 50.0 8.7 437 83.0
ConvNeXt-S [liu2022convnet] 2242 50 8.7 447 83.1
Iwin-S(ours) 2242 51.6 9.0 442/464 83.4
@@
DeiT-Base/16[touvron2020deit] 2242 86.6 17.6 273 81.8
T2T-ViT-24[T2T] 2242 64.1 14.1 - 82.3
PVT-Large[PVT_v1] 2242 61.4 9.8 357 81.7
Twins-L[chu2021twins] 2242 99.2 15.1 271 83.7
PVT-v2-B4[PVT_v2] 2242 62.6 10.1 298 83.6
PoolFormer-M48[metaformer] 2242 73.0 11.8 337 82.5
Swin-B[liu2021swin] 2242 88.0 15.4 287 83.3
Swin-B[liu2021swin] 3842 88.0 47.0 85 84.5
ConvNeXt-B [liu2022convnet] 2242 89 15.4 292 83.8
ConvNeXt-B [liu2022convnet] 3842 89 45.0 96 85.1
@@
R-101x3 [kolesnikov2020bigtransferbitgeneral] 3842 388 204.6 - 84.4
R-152x4 [kolesnikov2020bigtransferbitgeneral] 4802 937 840.5 - 85.4
ViT-B/16 [dosovitskiy2021imageworth16x16words] 3842 86 55.4 86 84.0
ViT-L/16 [dosovitskiy2021imageworth16x16words] 3842 307 190.7 27 85.2
Swin-B [liu2021swin] 2242 88 15.4 278 85.2
Swin-B [liu2021swin] 3842 88 47.0 85 86.4
ConvNeXt-B [liu2022convnet] 2242 89 15.4 292 85.8
ConvNeXt-B [liu2022convnet] 3842 89 45.1 96 86.8
@@
Swin-L [liu2021swin] 2242 197 34.5 145 86.3
Swin-L [liu2021swin] 3842 197 103.9 46 87.3
ConvNeXt-L [liu2022convnet] 2242 198 34.4 147 86.6
ConvNeXt-L [liu2022convnet] 3842 198 101.0 50 87.5
@@ IV-B Object Detection on COCO
We follow Swin [liu2021swin] settings and evaluate the Iwin backbone on the COCO [lin2014coco] for both object detection and instance segmentation tasks using Mask R-CNN [he2017mask] and Cascaded Mask R-CNN [cai2018cascade] with the MMDetection [mmdetection] toolbox. We use multi-scale training, AdamW optimizer and pre-trained models on ImageNet-1K.
@@ Results
[…]
@@
[…]
@@ IV-C Semantic Segmentation on ADE20K
We evaluate the Iwin backbone on the ADE20K [zhou2019semantic] semantic segmentation benchmark using UperNet [xiao2018upernet] with the MMSegmentation [mmseg2020] toolbox. The experimental settings follow Swin [liu2021swin]. Training is performed for 160K iterations with a total batch size of 16 (2 images per GPU across 8 GPUs).
@@ Results
[…]
@@
ResNet50[he2016deep] 28.5 183 36.7 - - -
PVTv2 B2[PVT_v1] 29.1 165 45.2 - - -
ConvNeXt-T[liu2022convnet] - - - 60.0 939 46.0
Swin-T[liu2021swin] - - - 59.9 945 44.5
@@
ResNet101[he2016deep] 47.5 260 38.8 - - -
PVTv2 B3[PVT_v1] 49.0 224 47.3 - - -
ConvNeXt-S[liu2022convnet] - - - 82.0 1027 48.7
Swin-S[liu2021swin] - - - 81.3 1038 47.6
@@
[…]
PVTv2 B4[PVT_v1] 66.3 285 47.9 - - -
ConvNeXt-B[liu2022convnet] - - - 122.0 1170 49.1
Swin-B[liu2021swin] - - - 121.0 1188 48.1
@@ IV-D Video Recognition on Kinetics-400
In line with the Video Swin Transformer [liu2021video] settings, we propose the Video Iwin Transformer for action recognition with the MMaction2 [2020mmaction2] toolbox on the Kinetics-400 [kay2017kinetics] dataset. The Video Iwin Transformer is initialized with its pre-trained Iwin model from ImageNet. Unlike the Video Swin Transformer’s intricate adaptations from 2D to 3D, which involve window shifting, masking, and relative positioning, the Iwin model only adjusts the window size and convolution kernel size and leaves the rest unchanged. As illustrated in Figure 9, in Iwin 3D Attention, tokens from interleaved windows across all frames are collected for attention calculation, while depthwise separable convolution maintains its 2D operation on individual frames.
@@ Results
Additionally, we explore the impact of window size on computational resources and performance. The evaluation 10 demonstrates a distinct trade-off between computational efficiency and predictive performance across different configurations of the $\mathrm{Iwin\text{-}T}$ architecture, governed by the variations in attention window volume ($W$, defined by window size $w$) and convolution kernel dimensions ($k$). As the spatiotemporal window volume expands from the compact $\mathrm{Iwin\text{-}T}_{\mathrm{w277}}^{\mathrm{k233}}$ ($W=98$) to the extensive $\mathrm{Iwin\text{-}T}_{\mathrm{w1677}}^{\mathrm{k133}}$ ($W=784$), we observe a monotonic increase in computational overhead, where memory usage and iteration latency escalate by approximately $3.2\times$ and $1.6\times$, respectively. While the enlargement of the temporal window in the $\mathrm{w1677}$ variant ($16\times 7\times 7$) effectively captures longer-range temporal dependencies to achieve the peak accuracy of $79.06\%$, this marginal accuracy gain comes at a disproportionate cost in FLOPs and structural complexity compared to the smaller $\mathrm{w277}$ variant ($78.22\%$). This analysis indicates that while the $\mathrm{Iwin\text{-}T}_{\mathrm{w1677}}^{\mathrm{k133}}$ configuration maximizes precision by leveraging an extensive receptive field, the models with moderate window sizes offer a more balanced efficiency-accuracy profile, minimizing resource consumption while maintaining competitive classification capability. Radar chart11 provides a more intuitive view of the differences between them.
Figure 9: Illustration of Iwin 3D Attention. In (a), depth-wise convolution is performed on each frame, as in 2D, and window attention is performed in 3D. Here, in addition to spatial position encoding and collaboration with window attention, the convolution plays a third role: telling the model which tokens are in the same frame, similar to frame attention introduced by VGGT [wang2025vggt], thus avoiding temporal position encoding, even though there is no RTR in the temporal dimension. In (b), we additionally perform the RTR operation in the time dimension and cooperate with a 3D depth-wise convolution, with its kernel size in the time dimension spanning frames. (a) can be regarded as a special case of (b). This effectively approximates 3D full attention, allowing connections between any tokens within a video. No position coding is required for the time dimension for both methods.
[…]
@@
R(2+1)D [tran2018closer] - 72.0 90.0 10 × 1 75 61.8
I3D [carreira2017quo] ImageNet-1K 72.1 90.3 - 108 25.0
NL I3D-101 [wang2018non] ImageNet-1K 77.7 93.3 10 × 3 359 61.8
SlowFast R101+NL [feichtenhofer2019slowfast] - 79.8 93.9 10 × 3 234 59.9
X3D-XXL [feichtenhofer2020x3d] - 80.4 94.6 10 × 3 144 20.3
MViT-B, 32×3 [fan2021multiscale] - 80.2 94.4 1 × 5 170 36.6
MViT-B, 64×3 [fan2021multiscale] - 81.2 95.1 3 × 3 455 36.6
ViViT-L/16x2 [arnab2021vivit] ImageNet-21K 80.6 94.7 4 × 3 1446 310.8
ViViT-L/16x2 320 [arnab2021vivit] ImageNet-21K 81.3 94.7 4 × 3 3992 310.8
Swin-T [liu2021swin] ImageNet-1K 78.8 93.6 4 × 3 88 28.2
Swin-S [liu2021swin] ImageNet-1K 80.6 94.5 4 × 3 166 49.8
[…]
Figure 10: Impact of Spatiotemporal Window Volume on Computational Resources and Performance. Larger window volume results in greater computational cost but higher accuracy.
Figure 11: Multidimensional Trade-off Analysis of Efficiency and Accuracy across Video Iwin-T Variants on the Kinetics-400.
@@ IV-E Image Generation
[…]
@@ Results
As shown in Table VI, our proposed FlashDiT demonstrates efficiency in image generation. It achieves competitive performance in only 56 training epoches, a fraction of the epoches required by previous state-of-the-art models such as DiT (1400) and MAR (800). Unlike these models have complexity of $32^{2}$ or $16^{2}$, FlashDiT has a computational complexity of only $3^{2}+4^{2}=25$ while maintaining high generation quality, achieving a gFID of 3.08 and an IS of 223.2 without classifier guidance. Therefore, FlashDiT verifies the effectiveness of the key component in Iwin as a standalone module that can seamlessly replace the self-attention module in generation models.
@@
Method Epoches Params $\times256$ w/o CFG $\times256$ w/ CFG
@@
LlamaGen [llamagen] 300 3.1B 9.38 8.24 112.9 0.69 0.67 2.18 5.97 263.3 0.81 0.58
VAR [var] 350 2.0B - - - - - 1.80 - 365.4 0.83 0.57
MagViT-v2 [magvitv2] 1080 307M 3.65 - 200.5 - - 1.78 - 319.4 - -
MAR [mar] 800 945M 2.35 - 227.8 0.79 0.62 1.55 - 303.7 0.81 0.62
@@
MaskDiT [maskdit] 1600 675M 5.69 10.34 177.9 0.74 0.60 2.28 5.67 276.6 0.80 0.61
DiT [peebles2023scalable] 1400 675M 9.62 6.85 121.5 0.67 0.67 2.27 4.60 278.2 0.83 0.57
SiT [sit] 1400 675M 8.61 6.32 131.7 0.68 0.67 2.06 4.50 270.3 0.82 0.59
MDTv2 [mdtv2] 1080 675M - - - - - 1.58 4.52 314.7 0.79 0.65
REPA [repa] 800 675M 5.90 - - - - 1.42 4.70 305.7 0.80 0.65
LightningDiT [yao2025vavae] 64 675M 5.14 4.22 130.2 0.76 0.62 2.11 4.16 252.3 0.81 0.58